# Experience Retrieval-Augmentation with Electronic Health Records Enables Accurate Discharge QA

### v1

@@ 5.3 Further Analysis
We conducted additional studies to investigate the impact of key components on the performance of ExpRAG, including the number of similar patients $k$, the balancing coefficients $\lambda_{1/2/3}$, and the choice of retrievers. We apply GPT-3.5 as backbone by default.
@@
Balacing Coefficients As shown in Equ. 4, we introduce three coefficients to balance the similarity computed based on diagnosis, medication, and procedures. We apply an equal distribution by default and explore the effect of using different weighting strategies here:
@@
Top-$k$ Patients We vary the number of retrieved similar patients $k$ on QA performance. As shown in Table 4, different tasks have varies trends. For example, on Instruction task, accuracy starts to increases with larger number, from $k=10$ with an accuracy of 78.75. While for Diagnosis and Medication tasks, beyond $k=20$, performance fluctuates, suggesting that while more retrieved candidates provide useful context, excessive retrieval may introduce irrelevant or conflicting information, leading to slight declines in accuracy.
@@
Table 4: Performance Comparison with Different $k$ for GPT3.5.
Experience Retriever We evaluate four retrievers for experience retrieval: auto-merging Liu (2022), sentence-window Liu (2022), BM25 Robertson et al. (2009), BM25+ (the combination of BM25 and word embeddings), Contriever Izacard et al. (2021), and flare Jiang et al. (2023b). More details regarding the methods can be found in Appendix A.2.
As shown in Table 5, auto-merging, sentence-window, and flare achieve the best performance, highlighting the effectiveness of context-aware retrieval. However, Contriever, which utilizes unsupervised dense representations, also underperforms in our cases, suggesting the need for medical domain-specific fine-tuning.
@@
5.4 Case Studies
@@
We conduct a case study focusing on Discharge Diagnosis in DischargeQA. We examine the discharge diagnosis of a target patient (ID: 20453584) who presented with bilateral ulnar paresthesias and neck pain. We compare this patient with similar patients, who are selected by comparing ICD/NDC codes from EHR tabular data. Fig 4 shows one example question in DischargeQA and the similarities in the discharge reports between the target patient and the similar patients with IDs 25633130, 29378221 and 28817667.
Figure 4: Comparison of Similar Patients.
Analysis of Similar Patients’ Data: Upon reviewing the discharge summaries of the similar patients, it became clear that several key diagnostic features were shared with patient 20453584:
@@
6 Conclusion
Inspired by the importance of experience in clinical decision-making, we propose a novel coarse-to-fine retrieval framework, ExpRAG, to utilize knowledge from similar patient records. Specifically, we introduce EHR as a knowledge basis and employ a comprehensive similarity measurement algorithm to narrow down the candidate pool, from which a text retriever extracts the relevant content. Evaluated on our curated DischargeQA, ExpRAG consistently improves the performance of various LLMs, highlighting the potential of leveraging past clinical experience for enhanced medical QA.
Limitations While EHR provides abundant medical information, such as lab test results, our proposed ExpRAG currently utilizes only diagnosis, medication, and procedures as an initial exploration. Additionally, DischargeQA currently consists solely of multi-option questions, which may limit its ability to comprehensively evaluate the generative capabilities of LLMs.
Appendix A Appendix
A.1 Dataset Details
A.2 Retriever
A.3 LLM Backbone
Deepseek-R1 DeepSeek-R1 is a powerful 671 billion parameter language model developed by DeepSeek AI, from which we use DeepSeek-R1-Distill-Llama-8B was derived as a more efficient 8 billion parameter version, distilled from the original model’s knowledge to offer improved performance on reasoning tasks while maintaining computational efficiency
A.4 Prompt Templates
Figure 6: Prompt design example for Diagnosis tasks

### v2

@@ 5.3 Further Analysis
We conducted additional studies to investigate the impact of key components on the performance of ExpRAG, including the number of similar patients $k$, and the balancing coefficients $\lambda_{1/2/3}$. GPT-3.5 is applied as the backbone by default.
@@
Balancing Coefficients. As shown in Equ. 4, we introduce three coefficients to balance the similarity computed based on diagnosis, medication, and procedures. We apply an equal distribution by default and explore the effect of using different weighting strategies here:
@@
Top-$k$ Patients. We vary the number of retrieved similar patients $k$ on QA performance. As shown in Table 6, different tasks have varies trends. For example, on Instruction task, accuracy starts to increases with larger number. While for Diagnosis and Medication tasks, beyond $k=20$, performance fluctuates, suggesting that while more retrieved candidates provide useful context, excessive retrieval may introduce irrelevant or conflicting information, leading to slight declines in accuracy.
@@
Table 6: GPT-3.5 performance with different $k$.
5.4 Case Study
We perform a focused case study on a Discharge Diagnosis task from DischargeQA. Specifically, we analyze a patient (ID: 20453584) presenting with bilateral ulnar paresthesias and neck pain. Similar patients are identified by matching ICD/NDC codes from structured EHR data (Figure 4). Reviewing discharge summaries of these similar patients revealed shared key diagnostic features—including cervical disc herniation, spinal stenosis, and upper extremity neurological symptoms—which substantially clarified the target patient’s clinical picture. For instance, similar patients exhibiting C6-C7 disc herniations and spinal stenosis provided critical evidence, improving the interpretation of the target patient’s symptoms and supporting a more accurate final diagnosis. We elaborate the explanation in Appendix A.8.
6 Conclusion
Inspired by the importance of experience in clinical decision-making, we propose a novel coarse-to-fine retrieval framework, ExpRAG, to utilize knowledge from similar patient records. Specifically, we introduce EHR as a knowledge basis and employ a reliable similarity measurement algorithm to narrow down the candidate pool to have relevant and useful content. Evaluated on our curated DischargeQA, ExpRAG consistently improves the performance of various LLMs, highlighting the potential of leveraging past experience to enhance model performance on medical QA.
Limitations While EHR provides abundant medical information, such as lab test results, our proposed ExpRAG currently utilizes only diagnosis, medication, and procedures as an initial exploration, which provides valuable insights and promising directions for future research. Additionally, DischargeQA currently consists solely of multi-option questions, which can be enhanced to be open-ended to comprehensively evaluate the generative capabilities of LLMs, When more economical and accurate LLMs are developed.
Appendix A Appendix
Figure 5: DischargeQA generation workflow.
A.1 DischargeQA
To evaluate LLMs in real-world clinical scenarios, we constructed a dataset combining structured tables from MIMIC-IV and unstructured discharge notes from MIMIC-IV-note, totaling over 140 000 patient records. The whole process is shown in 5
Patient selection.
As in  4.2, a valid pool of candidates for target or similar patients is the key point for retrieving information from similar patients.
Starting from the 430 k patients in MIMIC-IV, we retain (i) encounters with a discharge summary, (ii) 3–40 rows in each of diagnoses, prescriptions, and procedures, and (iii) discharge notes that contain $\geq 4$ instruction bullet points identified by GPT-4o. This yields 28 000 admissions, each can give enough structured items and narrative content to form challenging QA pairs, so that there are enough information to compare to other patients, and to have candidate options as wrong answer if there is no overlap in discharge reports.
Note segmentation and exposure.
Each discharge summary is heuristically split into seven sections (Patient Demography, Presenting Condition, Clinical Assessment, Treatment Plan, In-Hospital Progress, Discharge Summary, Post-Discharge Instructions) and mapped to three temporal phases: pre-diagnosis, in-hospital, and post-discharge, matching the 3 sections Clinical profile, In-hospital, and Discharge Plan as Figure  7. For every task we reveal only the phases that would have been available to the clinician at decision time, preventing label leakage.
Problem Design with EHR
  • Golden answers are clinician-authored discharge–note items.
  • Distractors are drawn from the same patient’s structured tables (drugs for medication, ICD-coded diagnoses for diagnosis).
  • Overlaps between structured candidates and golden answers are merged via GPT-4o.
  • Unlike single-choice formats, both diagnosis and medication tasks use multi-select, requiring selection of all correct items, thus raising task difficulty.
Construction Prompt Design
Prompts to generate options are carefully crafted to combine EHR tabular data and golden answer from discharge note. As an example, the prompt as Figure 8 defines the role of the model as a clinician and provides three key components: the list of discharge diagnoses, a database of historical diagnoses, and summarized background information extracted from the discharge summary. The task requires the model to identify which diagnoses should be included in the discharge summary by reasoning through the given data. Correct options are derived from the discharge diagnosis list, while incorrect but plausible options are generated from the diagnoses database. To ensure realism, GPT-4o is used to handle overlap, summarize long diagnoses, and align outputs with clinical expectations, providing a rigorous framework for evaluation.
A.2 Compare Two Rankers using similar patients
Figure 6: Text-Based Ranker vs EHR-Based Ranker workflow.
Shown as Figure  6.
A.2.1 Conventional RAG
A traditional RAG pipeline for clinical question answering consists of:
  1. 1.
    Encoding a query (e.g., a discharge planning question) into an embedding vector.
  2. 2.
    Chunking a large corpus into smaller text passages and embedding each chunk.
  3. 3.
    Finding top-k relevant text chunks by comparing similarity scores between query embeddings and chunk embeddings.
  4. 4.
    Retrieving the most relevant text passages.
  5. 5.
    Feeding the retrieved text to the LLM for answer generation.
However, applying this approach to EHR discharge summaries presents challenges: Inefficiency: The MIMIC-IV discharge summary corpus is over 4GB in size, making fine-grained chunk retrieval computationally expensive. Loss of Context: Traditional chunking disrupts the continuity of patient history, making it difficult for LLMs to infer longitudinal medical decisions. Scattered Information: Important information may be spread across multiple notes, making individual paragraph retrieval suboptimal.
A.2.2 Text-Based Ranker
Instead of retrieving individual paragraphs or sentences, we treat each patient’s discharge summary as a retrievable document and adapt the RAG pipeline accordingly:
Query Encoding Convert the question, options and target patient’s background into an embedding vector.
Patient-Level Indexing Store each patient’s discharge summary as a separate retrievable document and compute its embedding.
Find Similar Patients Using Embeddings Rank top-k similar patients from all patients based on similarity between the query embedding and patient embeddings.
Retrieve Top-K Patient Summaries Select the top-k most similar patient discharge summaries as a document source, like "experience".
Run RAG on Retrieved Summaries Use the retrieved summaries as references for LLM-based answer generation. Traditional RAG is running by embedding the query and the document source, retrieving chunks as reference, and answering query to generate the correct answer.
By preserving full discharge summaries, Text-based retrieval maintains patient-level coherence while improving retrieval efficiency.
A.2.3 EHR-Based Ranker
To further improve retrieval relevance, we introduce an EHR-based Ranker that utilizes structured patient data for retrieving better discharge summaries as the "experience" document source:
Identify the Target Patient’s Condition Extract structured EHR tabular data (ICD codes for diagnoses, prescriptions, and procedures) from the target patient.
Find Similar Patients Using Structured EHR Data Compute similarity between the target patient’s structured data and all other patients based on shared ICD codes in MIMIC-IV, after scanning all recorded patient data. Jaccard index is used to calculate the similarity.
Retrieve Top-K Similar Patients Select the top-k most similar patients based on their structured medical records.
Extract Their Discharge Summaries Retrieve the corresponding discharge summaries for the top-k similar patients.
Run RAG on Retrieved Summaries Feed the retrieved summaries into an LLM for answer generation, same as the last step in Text-Based Ranker.
By incorporating structured EHR data, this method ensures that retrieval is clinically relevant, going beyond semantic text similarity. This process resembles an experience-based search: identifying similar past experiences and adapting them to solve the problem at hand.
A.3 Dataset Details
A.4 Retriever
@@
A.4.1 Retriever Experiments
Our experiments include: Auto-merging Liu (2022), sentence-window Liu (2022), BM25 Robertson et al. (2009), BM25+ (the combination of BM25 and word embeddings), Contriever Izacard et al. (2021), and flare Jiang et al. (2023b). More details regarding the methods can be found in Appendix A.4. As shown in Table 7, auto-merging, sentence-window, and flare achieve the best performance, highlighting the effectiveness of context-aware retrieval. However, Contriever, which utilizes unsupervised dense representations, also underperforms in our cases, suggesting the need for medical domain-specific fine-tuning.
A.5 LLM Backbone
A.5.1 Proprietary Models
We selected GPT-3.5 and GPT-4o as representative models due to their extensive real-world adoption and practical relevance demonstrate the applicability of our framework in everyday clinical settings.
A.5.2 Open-source Models
Qwen3, Deepseek-R1 and Mistral were included as leading open-source models to benchmark the generalizability and robustness of our approach. Specifically, Qwen3 and Deepseek-R1 are thinking models, which is an essential feature to tackle challenges in our tasks.
Deepseek-R1-8B DeepSeek-R1 is a powerful 671 billion parameter language model developed by DeepSeek AI, from which we use DeepSeek-R1-Distill-Llama-8B was derived as a more efficient 8 billion parameter version, distilled from the original model’s knowledge to offer improved performance on reasoning tasks while maintaining computational efficiency
Qwen3-30B-A3B Qwen3-30B-A3B is developed by Alibaba Cloud, released at April 2025. It is a dense-and-MoE model that has a “thinking mode” for deep reasoning, math, and coding. It outperforms earlier Qwen generations on logical reasoning, code generation, tool-using agent tasks, and human-preference benchmarks, while supporting 100 + languages for instruction following and translation. In our work we leverage its thinking mode to maximize accuracy on complex reasoning tasks.
A.5.3 Medical Models
Baichuan-M1 Trained from scratch on 20 T tokens that mix high-quality clinical and general texts, Baichuan-14B-M1 is the first open-source 14 B-parameter LLM purpose-built for medicine. It incorporates specialised heads, enabling fine-grained reasoning that matches general-domain peers on standard benchmarks yet surpasses models 5× larger on medical tasks. An updated architecture with longer-context handling further improves comprehension of lengthy clinical narratives and complex patient histories.
We have experiments shown in  A.5.3. Baichuan-14B-M1 exhibits the same limitations we observed in other compact models. When additional context from EHR- or text-based retrieval is supplied, the model occasionally confuses the target patient with the retrieved similar cases. Under RAG settings, this confusion produces > 5 % invalid responses—outputs that either fail our answer-parsing patterns or omit a decisive answer—leading to a noticeable overall drop in accuracy. Despite these invalid cases, Baichuan-M1 using EHR-Based ExpRAG still surpasses other approaches on the Medication task, underscoring its strength in pharmacological reasoning even when other aspects falter.
Model Context Instruction Diagnosis Medication
Acc(%) F1 F1
Baichuan Direct-Ask 89.8 0.381 0.256
Text-based 87.8 0.377 0.277
ExpRAG ${}_{\text{EHR}}$ 86.3 0.373 0.285
Table 8: Performance from Biomedical Model, with 5% invalid answer on Text-based and EHR-based
UltraMedical Llama-3-8B-UltraMedical is derived from Meta’s Llama-3-8B and further tuned on the 410 K-example UltraMedical dataset (synthetic + curated), an 8 B-parameter biomedical specialist. It tops MedQA, MedMCQA, PubMedQA and MMLU-Medical, handily beating larger general models such as GPT-3.5 and Meditron-70B, and rivals domain-tuned Flan-PaLM and OpenBioLM-8B. The model targets exam-style question answering, literature comprehension and clinical-knowledge retrieval, making advanced medical NLP accessible at modest compute cost.
UltraMedical’s 70 B variant has a 8 k-token context window. When we integrated the model into our retrieval-augmented generation pipeline, more than 50 % of the assembled prompts—target patient record + similar-patient excerpts + task instruction—were longer than 8 k tokens. These over-length inputs had to be truncated or rejected, which systematically stripped clinical details from half of our queries and degraded answer quality. Because this hard context limit blocked reliable end-to-end evaluation, we discontinued UltraMedical-70B for our study despite its otherwise strong domain results.
A.6 Prompt Templates
A.7 Detailed Setup for Ranker Evaluation
Patient and Candidate Sampling.
We first randomly select 100 target patients from the MIMIC-IV cohort. For each target, we construct 100 candidate patients by:
  • Sampling 20 uniformly at random from the filtered step described in 4.2 from full dataset (to include unrelated cases).
  • Sampling 80 from a restricted pool formed by taking the patients (per target) with non-zero EHR similarity in either of three modalities (diagnoses, procedures, prescriptions).
Annotation with GPT4o-mini.
Each target–candidate pair is judged by GPT4o-mini along the three modalities. We average these three scores to obtain a single “ground truth” similarity for correlation.
Ranking Methods.
  1. 1.
    Text embeddings: Cosine similarity over discharge summaries using three pretrained models (bge-small-en-v1.5, all-MiniLM-L6-v2, paraphrase-MiniLM-L3-v2).
  2. 2.
    EHR-based EXPRAG: Jaccard similarity on code sets (ICD/NDC) for each modality, aggregated with equal weights.
Correlation Metrics.
For each ranker and target patient, we produce a ranked list of 100 candidates. We then compute Pearson’s $\rho$ and Spearman’s $\tau$ between the ranker’s scores and the GPT4o-mini annotations, and average each metric across all 100 targets.
Results.
Table 4 (main text) reports the final average correlations, demonstrating that EHR-based EXPRAG best aligns with human-like judgments.
Figure 8: Prompt design example for Diagnosis tasks
A.8 Details in Case Studies
@@
We conduct a case study focusing on Discharge Diagnosis in DischargeQA. We examine the discharge diagnosis of a target patient (ID: 20453584) who presented with bilateral ulnar paresthesias and neck pain. We compare this patient with similar patients, who are selected by comparing ICD/NDC codes from EHR tabular data. Fig4 shows one example question in DischargeQA and the similarities in the discharge reports between the target patient and the similar patients with IDs 25633130, 29378221 and 28817667. Upon reviewing the discharge summaries of the similar patients, it became clear that several key diagnostic features were shared with patient 20453584: